OpenStudy (anonymous):

suppose z=f(x,y) is a differentiable function on R^2 and x=rcos(theta) and y=rsin(theta). compute the partial of z with respect to r.

OpenStudy (dumbcow):

what is f(x,y)?

OpenStudy (anonymous):

according to the question it can be anything.

OpenStudy (anonymous):

damn. is this calc 2?

OpenStudy (dumbcow):

dx/dr = cos(theta) dy/dr = sin(theta)

OpenStudy (anonymous):

calc 3

OpenStudy (dumbcow):

hmm say z = x+y then dz/dr = cos(theta) + sin(theta) i think, dang i have forgotten a lot of calc 3

OpenStudy (anonymous):

lol. im in calc 1 :D cant rly help you

OpenStudy (dumbcow):

\[\frac{dz}{dr} = \frac{dz}{dx}\frac{dx}{dr} + \frac{dz}{dy}\frac{dy}{dr}\]

OpenStudy (anonymous):

that's what I was thinking. that's the answer i have. do you have a link or something to that? i'll need that again for future reference.

OpenStudy (dumbcow):

no that i actually remembered i tried looking it up but no luck yet...

OpenStudy (anonymous):

i know i've seen that somewhere myself. it was a vague memory but since you thought of the same thing i feel a lot better about it.

OpenStudy (anonymous):

I agree with you guys it's called the total derivative. http://en.wikipedia.org/wiki/Total_derivative. The case you are interested in is a little bit down the page

OpenStudy (anonymous):

thanks for the link. i feel much better now that two others agree.

OpenStudy (dumbcow):

thank you

OpenStudy (anonymous):

can either of you help with lagrange multipliers?

OpenStudy (anonymous):

for some reason my calc teachers never mentioned them.

OpenStudy (anonymous):

what's the problem though I'll try to look it up and maybe I can help?

OpenStudy (anonymous):

just posted it